# appendices

### v2

@@ Methods
Figure 4: Overview of the 18 areas analyzed by UNOSAT in Ukraine.
@@
Reference Damage Assessments
@@
To the best of our knowledge, the only public, georeferenced datasets that provide building-level damage information for Ukraine are the UNOSAT maps [12]. They are based on commercial VHR imagery, where analysts manually compare pre- and post-event images using a standardized damage scale. To maximize the chances that these damages produce discernible changes in the SAR imagery, we only keep the two most severe damage levels, i.e. destroyed or severely damaged, and discard the others. We aggregate all available UNOSAT maps for Ukraine into a dataset consisting of 10,934 unique entries, distributed across 18 different areas of interest (AOIs) as depicted in Fig. 4. Importantly, these labels are attached to geo-coordinates, without any relation to concrete events. We, therefore, lack information on both the spatial extent and the precise date of the destruction, all that is known is that it happened between the pre- and post-event dates. For the present work, we further restrict this interval by assuming that no damage occurred before the onset of the invasion on February 24, 2022. We point out that this assumption may not strictly hold in the eastern regions, particularly Luhansk and Donetsk oblasts. A detailed summary of our dataset can be found in Appendix A.
SAR Image Time Series
@@
We use freely available SAR data from the Copernicus Sentinel-1 mission [39], consisting of two satellites equipped with C-band SAR instruments and moving in the same orbital plane, phased 180° apart. This setup originally allowed for global coverage with a revisit period below 6 days anywhere on Earth. Unfortunately, this period has doubled to up to 12 days since the technical failure of Sentinel-1B in December 2021. For our study, we only use data from Sentinel-1A to maintain consistent temporal resolution. We utilize the Ground Range Detected (GRD) product available on GEE, consisting of log-amplitudes for the VV and VH polarizations, resampled at 10 m. Internally, GEE automatically preprocesses every tile with precise orbit correction, border and thermal noise removal, radiometric calibration, and terrain correction. We do not add any further preprocessing.
Building Footprints
@@
We leverage the building layer from the Overture Maps Foundation’s Open Map Data [45], which combines building footprints from diverse sources. In Ukraine, this equates to a total of 26.5M buildings, with 24.7% coming from OpenStreetMap’s crowd-sourced database [50] and the remaining 20M from Microsoft’s Global Building Footprints [51]. We note here that these building footprints have their own limitations. E.g., OpenStreetMap data may be outdated, and Microsoft footprints, retrieved from VHR optical imagery with a deep learning model, may suffer from missing buildings and geometric inaccuracies, particularly for small buildings. Considering the resolution of Sentinel-1 we only take into account the 15M buildings that exceed 50 m2 in surface area, i.e., they are larger than half a pixel. Appendix E illustrates some of these limitations.
Definition of Time Intervals
For clarity, we first define the temporal acquisition windows used in our study. Time interval $T_{0}$ covers one year from Feb. 24, 2020 until Feb. 23, 2021. The subsequent intervals $T_{1}$ to $T_{12}$ represent consecutive 3-month time windows, with $T_{1}$ ranging from Feb. 24, 2021 to May 23, 2021, and the last interval $T_{12}$ spanning Nov. 24, 2023 to Feb. 23, 2024. As such, intervals $T_{1}$ to $T_{4}$ represent the year preceding the invasion, while $T_{5}$ to $T_{12}$ represent the two years following it.
@@
Figure 5: Overview of our machine learning framework. For training (left), we use per-pixel Sentinel-1 time series extracted at the location of UNOSAT point annotations. The model is fed with a pair of time series from the same location. The first one spans a fixed 12-month time interval $T_{0}$ from 2020, and the second one spans one of the 3-month time intervals $T_{n}$ between 2021 and 2023. Both time series are encoded with a custom features extractor, and damage labels are dynamically assigned according to $T_{n}$. At inference time (center), the model generates a damage probability heatmap valid at $T_{n}$ and spanning the entire country, aggregating the predictions of different Sentinel-1 orbits. The raw damage probabilities are intersected with building footprints. For the final map the estimates for different time intervals $T_{n}$ are thresholded and aggregated.
Time Series Extraction
@@
For each location marked by UNOSAT, we stack the corresponding Sentinel-1 tiles and extract all backscatter values, for both the VV and VH polarizations. Limiting the training to the points annotated by UNOSAT corresponds to a fully supervised learning scheme where the target value is known for every training example. Relying only on pixel-wise time series discards all spatial context. We found that the time series signal carries most of the information relevant to our analysis, presumably because conflict-induced building damages are mostly small and localized. Per-pixel processing also scales particularly well on parallel processing architectures like the one underlying GEE.
@@
Importantly, Sentinel-1 has the typical side-looking viewing geometry of SAR sensors, where the same location may look very different depending on the incidence angle and orbit direction. We therefore extract the backscatter values independently for each orbit, resulting in 2 to 4 time series per UNOSAT reference point. Overall, we obtain 33,304 distinct time series, denoted as $x_{i}^{o}$, where $i$ denotes the ID (respectively, location) and $o$ is the orbit index.
Classification
@@
We formulate damage mapping as a supervised, binary classification. From each time series, we extract two segments: the fixed reference period $x_{i,\text{ref}}^{o}$ and the assessment period $x_{i,\text{new}}^{o}$. The task of the classifier is to estimate the probability that a war-related change has occurred in the time window between the two segments. For each segment and each band, we extract a set of $N$ statistical features and combine them into a final feature vector denoted as $\phi(x_{i})\in\mathbb{R}^{4N}$ that forms the input to the classifier. This strategy has several advantages. First, by fixing the reference period before the start of the war one can ensure that it represents the state without any war-induced damages. Second, extracting fixed-length segments makes the method independent of the overall conflict duration. Third, using multiple different assessment periods greatly increases the number of training examples and also helps to make the model robust against seasonal variations.
@@
As classification algorithm, we use a Random Forest as implemented in the SMILE library [52], since that algorithm is available off-the-shelf in GEE, facilitating large-scale deployment. For every input feature vector the model outputs a score $\hat{y}_{i}^{o}\in[0,1]$, which can be interpreted as the likelihood that war-induced damage has happened between the associated reference and assessment periods. At inference time, we compute the overall damage probability map by averaging the estimates from different orbits, $\hat{y}_{i}=\frac{1}{N_{o}}\sum\hat{y}_{i}^{o}$.
Training Details
@@
We perform all computations directly in GEE. We train the model on the four AOIs with the highest numbers of reference labels, Mariupol, NW Kyiv, Rubizhne, and Volnovakha. Together they represent 75.0% of all UNOSAT annotations and 82.6% of all backscatter time series. The remaining fourteen AOIs are reserved for evaluation, ensuring geographic diversity (see Fig. 4). We choose $T_{0}$ as a fixed reference period and randomly use periods from $T_{1}$ to $T_{8}$ (e.g. spanning one year before and after the beginning of the war) as assessment periods.
@@
To quantify the impact of the war on the Ukrainian building stock, we cross-reference our maps with the building footprints sourced from Overture Maps [45]. For each building $b_{j}$ and each period $T_{n}$, we assign a damage likelihood $\hat{y}_{j,T_{n}}$ by averaging the likelihood values of all pixels that fully or partially overlap the footprint, weighted by the overlap fraction:
@@
All the code necessary to reproduce our results can be found in the following repository: https://github.com/olidietrich/ukraine-damage-mapping-tool/.
@@
All data used in this study is publicly accessible. Links to all results, including damage heatmaps and building footprints with corresponding damage estimates, will be made available in the repository hosting our code.
Appendix A UNOSAT Dataset
Our reference dataset has been created by combining all existing UNOSAT datasets for Ukraine. We combined them into 18 distinct areas of interest (AOIs). For our purposes, we only use the two strongest labels, destroyed and severely damaged. For the sake of completeness, we also report the third one here, moderately damaged. For Mariupol and Kharkiv, for which multiple analyses were conducted by UNOSAT, we keep only the latest label assigned to a point, assuming that in the case of repeated assessments, the annotations were progressively refined. Table 4 summarizes our AOIs and the number of labeled locations in each of them.
@@ Appendix A UNOSAT Dataset
Table 4: Summary of the 18 distinct areas of interest (AOI) in Ukraine, merging data from 29 UNOSAT products. The table lists counts of destroyed, severely damaged, and moderately damaged labels per area, totaling 16,548 labeled damage instances, or 10,934 after excluding the moderately damaged ones.
Appendix B Detailed Comparison with the PWTT Baseline
The Pixel-Wise T-Test (PWTT), proposed in [43], is a statistical approach designed to identify changes by comparing the mean values of two Sentinel-1 amplitude time series, adjusted by their standard deviation. We use it as a performance baseline since it utilizes the same data and has similar specifications, in particular, the scalability to large regions and the ease of deployment on GEE.
@@ Appendix B Detailed Comparison with the PWTT Baseline
For a meaningful comparison, we keep the recommended, optimal cutoff value but rerun PWTT using our time periods and our own, stricter definition of negative labels. We report the corresponding results in Table 5. For readability, we repeat our results in Table 6. It should be noted that Chernihiv and Kharkiv, which represent 61% of our test set, were used by the original authors during the calibration of the PWTT. Nevertheless, our method consistently outperforms the baseline across all metrics.
@@
Table 5: Quantitative performance of the PWTT approach on our test set, with the cutoff value of 1.63 as recommended for Ukraine. All metrics are given as percentages.
@@
Table 6: Quantitative performance of our approach on our test set, with confidence threshold 0.5 (reprinted from Table 1). All metrics are given as percentages.
Appendix C Choice of threshold
@@
There is an inherent trade-off between precision and recall, and the optimal threshold depends on the specific application. To keep false alarms at a low enough rate, we set a target precision of 90%, for which we find an optimal confidence threshold of 0.655. That value has been used to generate all metrics in the present paper unless stated otherwise. With this threshold, the user can expect a recall of 46.5%. In practice, users of our tool can dynamically adjust the confidence threshold to suit their specific needs. For instance, one could start with a high threshold to discover hotpots with a very high likelihood of destruction, then gradually decrease the threshold to increase the recall and better assess the extent of destruction. Fig. 6 illustrates how our metrics vary with different thresholds.
@@
Figure 6: Pixel-level performance of our method on our test set, for different confidence thresholds. We set a target desired precision of 90% and found a confidence threshold of 0.655. The two vertical lines indicate the thresholds 0.5 and 0.655.
Appendix D Temporal Evolution of Damage
@@
The series of maps in Figure Fig. 7 illustrates the progression of building damage over the first two years of the conflict. The maps clearly delineate the primary locations of battles and the frontline, highlighting the areas most affected as the conflict persists.
@@
Figure 7: Percentage of buildings likely damaged for each period, cumulated since the beginning of the war and aggregated by hromadas. The predictions were thresholded at 0.655, and only buildings larger than 50 m2 were considered.
Appendix E Limitations of Building Footprints
@@
The building footprints provided by Overture Maps, which for Ukraine were compiled from OpenStreetMap and Microsoft, inevitably come with limitations. For instance, OpenStreetMap data may be obsolete due to the demolition of older structures (Fig. 8.A) or the construction of new ones (Fig. 8.B). While Microsoft footprints serve as a valuable complement to OpenStreetMap data, they were automatically retrieved with a deep neural network from Bing Maps imagery, which may not always offer recent images, and suffers from gaps due to cloudy or outdated acquisitions, for example in Kherson (Fig. 8.C). Moreover, it is not uncommon for large non-building structures, e.g. containers (Fig. 8.D), boats (Fig. 8.E) or even aircraft (Fig. 8.F), to be wrongly identified as buildings.
@@
Figure 8: Quality issues of existing building footprints. All images cover 500$\times$500 m2 and are displayed alongside the corresponding building footprints sourced from Overture Maps. Only buildings larger than 50 m2 are depicted. (A) Outdated buildings in Sevastopol. (B) Absent shopping mall near Chernihiv. (C) Missing footprints in Kherson. (D) Containers in Odessa. (E) Boats in Zaporizhzhia. (F) Antonov International Airport, Hostomel. Satellite images were obtained from ESRI (A and C), Bing (B, E, and F), and Google (D).
Appendix F World Bank Estimates
@@
In this section, we compare our estimates, with a total of $\approx$400,000 buildings damaged between Feb 2022 to Feb 2024, to figures provided by the World Bank (WB) in their third Rapid Damage and Needs Assessment report [2, p. 74-80].
@@
Fig. 9 shows the differences between our estimates and those of the WB. Excluding the other oblasts outside the direct conflict zone, both distributions align quite well, with indeed the majority of damage concentrated in Donetska and Kharkivska. However, our model estimates significantly more damage than the WB for the other oblasts, i.e. the ones further from the frontline. It appears that construction and other non-conflict-related changes dominate in those regions, which our model is unable to distinguish from building damages. This highlights that there is a trade-off between, on the one hand, comprehensive screening to detect also unexpected conflict-induced destruction far from the known frontline; and, on the other hand, accurate assessment by ruling out areas that are a-priori unlikely to be affected.
@@
Figure 9: Comparison of the geographical distribution of the World Bank estimates against ours. The World Bank provides spatial distribution for their damage cost estimates, while our spatial distribution shows buildings likely damaged.
Appendix G Changes in Backscatter due to Destruction
@@
The backscatter signal from a building is influenced by numerous variables, including building characteristics (e.g., geometry, orientation, surface texture, material composition), satellite parameters (e.g., frequency, polarization, incidence angle, orbit direction), and environmental factors (e.g., vegetation, soil moisture, snow, or heavy rainfall). Intuitively, one might expect that intact structures yield stronger specular returns, with flat surfaces and right-angle corners acting as corner reflectors, whereas structural damage would disrupt these patterns, producing a more diffuse signal. However, several studies indicate that damage signatures in SAR data are highly variable and depend on the type and extent of destruction [19, 20]. Damaged buildings can even exhibit backscatter signals as high as undamaged structures, or even higher [53].
@@
Our experiments corroborate these findings. Fig. 10 presents pixel-wise backscatter time series that illustrate how conflict-related destruction appears in the data. High-resolution pre- and post-destruction satellite imagery from Google Earth is provided for reference. We observe that the backscatter amplitude following destruction can vary widely – increasing, decreasing, or remaining largely stable – depending on the location, the satellite orbit, and the polarization. These examples emphasize the need to utilize all available orbits in order to maximize reliability.
@@
Figure 10: Sentinel-1 backscatter time series at selected UNOSAT locations for VV and VH polarizations, with VHR satellite images from Google Earth showing pre- and post-destruction conditions. Vertical dashed lines indicate the date of Russia’s full-scale invasion (black) and the dates of the displayed pre-destruction (green) and post-destruction (red) images. For readability, only two orbits per location are shown, even if in some locations up to four orbits are available. Amplitude can decrease (B, C), increase (A, C) or remain relatively stable (B, D) after destruction. The same event may manifest differently depending on the flight direction (B), or even on the incidence angle within the same flight direction (C).
Appendix H Ablation studies
@@
The $F_{1}$-score of our optimal configuration, which utilizes both VV and VH Sentinel-1 bands, all seven features, 50 decision trees and extracts the time series pixel-wise, is 0.72. Fig. 11 summarized the outcome for different settings. We observe the following characteristics:
@@
Input data: We tried replacing Sentinel-1 SAR data with Sentinel-2 optical images [54]. Pixel-wise Sentinel-2 time series were extracted from the Level-2A collection available on GEE [55], from which we filter out cloudy pixels using GEE’s CloudScore+ mask [56]. Various band combinations were tried, but incorporating Sentinel-2 data never enhanced model performance. This was expected, as our pipeline was design for Sentinel-1 properties, such as stable pre-event signals and a focus on temporal context over spatial context. Intuitively, it is harder to distinguish intact buildings from damaged ones by looking at the brightness change at an individual pixel location. Also combining Sentinel-1 and Sentinel-2 led to poorer results than using Sentinel-1 alone, likely due to the confounding effect of a six-fold increase in features, most of which contain little information for the task.
Model: We evaluated alternative classifiers available in GEE, i.e., Boosted Trees [52] and SVM [57]. Boosted trees showed weaker performance, while taking approximately 3$\times$ longer to run. We were unable to obtain results for the SVM: when applied to our data it quickly exceeded the computational budget allowed by GEE.

### v3

@@ Methods
Figure 4: Regions analyzed by UNOSAT. Overview of the 18 regions assessed by UNOSAT in Ukraine. Blue and orange dots indicate cities used for training and testing, respectively. The three insets highlight the variation in damage density across different regions. Background map: OpenStreetMap.
@@
All the data used in this work can be found freely available online.
@@
Reference Damage Assessments: To the best of our knowledge, the only public, georeferenced datasets that provide building-level damage information for Ukraine are the UNOSAT maps [unosat]. They are based on commercial VHR imagery, where analysts manually compare pre- and post-event images using a standardized damage scale. To maximize the chances that these damages produce discernible changes in the SAR imagery, we only keep the two most severe damage levels, i.e., destroyed or severely damaged, and discard the others. We aggregate all available UNOSAT maps for Ukraine into a dataset consisting of 10,934 unique entries, distributed across 18 different areas of interest (AOIs) as depicted in Fig. 4. Importantly, these labels are attached to geo-coordinates, without any relation to concrete events. We, therefore, lack information on both the spatial extent and the precise date of the destruction, all that is known is that it happened between the pre- and post-event dates. For the present work, we further restrict this interval by assuming that no damage occurred before the onset of the invasion on February 24, 2022. We point out that this assumption may not strictly hold in the eastern regions, particularly Luhansk and Donetsk oblasts. A detailed summary of our dataset can be found in Supplementary Note 1.
@@
SAR Image Time Series: We use freely available SAR data from the Copernicus Sentinel-1 mission [sentinel_1], consisting of two satellites equipped with C-band SAR instruments and moving in the same orbital plane, phased 180° apart. This setup originally allowed for global coverage with a revisit period below 6 days anywhere on Earth. Unfortunately, this period has doubled to up to 12 days since the technical failure of Sentinel-1B in December 2021. For our study, we only use data from Sentinel-1A to maintain consistent temporal resolution. We utilize the Ground Range Detected (GRD) product available on GEE, consisting of log-amplitudes for the VV and VH polarizations, resampled at 10 m. Internally, GEE automatically preprocesses every tile with precise orbit correction, border and thermal noise removal, radiometric calibration, and terrain correction. We do not add any further preprocessing.
@@
Building Footprints: We leverage the building layer from the Overture Maps Foundation’s Open Map Data [OvertureMaps], which combines building footprints from diverse sources. In Ukraine, this equates to a total of 26.5M buildings, with 24.7% coming from OpenStreetMap’s crowd-sourced database [OpenStreetMap] and the remaining 20M from Microsoft’s Global Building Footprints [MSFTfootprints]. We note here that these building footprints have their own limitations. E.g., OpenStreetMap data may be outdated, and Microsoft footprints, retrieved from VHR optical imagery with a deep learning model, may suffer from missing buildings and geometric inaccuracies, particularly for small buildings. Considering the resolution of Sentinel-1 we only take into account the 15M buildings that exceed 50 m2 in surface area, i.e., they are larger than half a pixel. Supplementary Note 5 illustrates some of these limitations.
@@
An overview of our machine learning framework is shown on Fig. 5. For a detailed ablation study of our key design choices, see Supplementary Note 8.
@@
Figure 5: Overview of the machine learning framework. For training (left), we use per-pixel Sentinel-1 time series extracted at the location of UNOSAT point annotations. The model is fed with a pair of time series from the same location. The first one spans a fixed 12-month time interval $T_{0}$ from 2020, and the second one spans one of the 3-month time intervals $T_{n}$ between 2021 and 2023. Both time series are encoded with a custom features extractor, and damage labels are dynamically assigned according to $T_{n}$. At inference time (center), the model generates a damage probability heatmap valid at $T_{n}$ and spanning the entire country, aggregating the predictions of different Sentinel-1 orbits. The raw damage probabilities are intersected with building footprints. For the final map the estimates for different time intervals $T_{n}$ are thresholded and aggregated.
Definition of Time Intervals: For clarity, we first define the temporal acquisition windows used in our study. Time interval $T_{0}$ covers one year from Feb. 24, 2020 until Feb. 23, 2021. The subsequent intervals $T_{1}$ to $T_{12}$ represent consecutive 3-month time windows, with $T_{1}$ ranging from Feb. 24, 2021 to May 23, 2021, and the last interval $T_{12}$ spanning Nov. 24, 2023 to Feb. 23, 2024. As such, intervals $T_{1}$ to $T_{4}$ represent the year preceding the invasion, while $T_{5}$ to $T_{12}$ represent the two years following it.
@@
Time Series Extraction: For each location marked by UNOSAT, we stack the corresponding Sentinel-1 tiles and extract all backscatter values, for both the VV and VH polarizations. Limiting the training to the points annotated by UNOSAT corresponds to a fully supervised learning scheme where the target value is known for every training example. Relying only on pixel-wise time series discards all spatial context. We found that the time series signal carries most of the information relevant to our analysis, presumably because conflict-induced building damages are mostly small and localized. Per-pixel processing also scales particularly well on parallel processing architectures like the one underlying GEE.
@@
Importantly, Sentinel-1 has the typical side-looking viewing geometry of SAR sensors, where the same location may look very different depending on the incidence angle and orbit direction (see Supplementary Note 7 for examples). We therefore extract the backscatter values independently for each orbit, resulting in 2 to 4 time series per UNOSAT reference point. Overall, we obtain 33,304 distinct time series, denoted as $x_{i}^{o}$, where $i$ denotes the ID (respectively, location) and $o$ is the orbit index.
@@
Classification: We formulate damage mapping as a supervised, binary classification. From each time series, we extract two segments: the fixed reference period $x_{i,\text{ref}}^{o}$ and the assessment period $x_{i,\text{new}}^{o}$. The task of the classifier is to estimate the probability that a war-related change has occurred in the time window between the two segments. For each segment and each band, we extract a set of $N$ statistical features and combine them into a final feature vector denoted as $\phi(x_{i})\in\mathbb{R}^{4N}$ that forms the input to the classifier. This strategy has several advantages. First, by fixing the reference period before the start of the war one can ensure that it represents the state without any war-induced damages. Second, extracting fixed-length segments makes the method independent of the overall conflict duration. Third, using multiple different assessment periods greatly increases the number of training examples and also helps to make the model robust against seasonal variations.
@@
As classification algorithm, we use a Random Forest as implemented in the SMILE library [Li2014Smile], since that algorithm is available off-the-shelf in GEE, facilitating large-scale deployment. For every input feature vector the model outputs a score $\hat{y}_{i}^{o}\in[0,1]$, which can be interpreted as the likelihood that war-induced damage has happened between the associated reference and assessment periods. At inference time, we compute the overall damage probability map by averaging the estimates from different orbits, $\hat{y}_{i}=\frac{1}{N_{o}}\sum\hat{y}_{i}^{o}$.
@@
Training Details: We perform all computations directly in GEE. We train the model on the four AOIs with the highest numbers of reference labels, Mariupol, NW Kyiv, Rubizhne, and Volnovakha. Together they represent 75.0% of all UNOSAT annotations and 82.6% of all backscatter time series. The remaining fourteen AOIs are reserved for evaluation, ensuring geographic diversity (see Fig. 4). We choose $T_{0}$ as a fixed reference period and randomly use periods from $T_{1}$ to $T_{8}$ (e.g., spanning one year before and after the beginning of the war) as assessment periods.
@@
To quantify the impact of the war on the Ukrainian building stock, we cross-reference our maps with the building footprints sourced from Overture Maps [OvertureMaps]. For each building $b_{j}$ and each period $T_{n}$, we assign a damage likelihood $\hat{y}_{j,T_{n}}$ by averaging the likelihood values of all pixels that fully or partially overlap the footprint, weighted by the overlap fraction:
@@
All the code necessary to reproduce our results can be found in the following repository: https://github.com/prs-eth/ukraine-damage-mapping-tool/.
@@
All data used in this study is publicly accessible. Links to all results, including damage heatmaps and building footprints with corresponding damage estimates, can be found in the repository hosting our code and/or in Zenodo (https://zenodo.org/records/14811504/).
Author Contributions
Olivier Dietrich: conceptualization (equal), data curation (lead), investigation (lead), methodology (lead), software (lead), visualization (lead), writing - original draft preparation (lead). Torben Peters: conceptualization (support), investigation (support), methodology (equal), software (support), supervision (equal), writing - original draft (equal). Vivien Sainte Fare Garnot: conceptualization (support), supervision (support), writing - review & editing (support). Valerie Sticher: conceptualization (lead), data curation (support), funding acquisition (lead), project administration (equal), writing - original draft (equal). Thao Ton-That Whelan: funding acquisition (equal) project administration (equal), validation (support), writing - review & editing (support). Konrad Schindler: conceptualization (equal), funding acquisition (equal), methodology (equal), project administration (equal), resources (lead), supervision (lead), writing - review & editing (lead). Jan Dirk Wegner: conceptualization (lead), funding acquisition (lead), methodology (equal), project administration (equal), resources (support), supervision (lead), writing - review & editing (lead).
Competing Interests
The authors declare no competing interests.
Acknowledgments
This work was supported in part by ETH4D’s Engineering Humanitarian Action initiative through the Remote Monitoring of Armed Conflict project in collaboration with the ICRC, as well as by ESA’s Discovery Programme under the Humanitarian Monitoring with Copernicus Satellites project (Contract No. 4000138343).
[…]
[title=References]
Supplementary Material
[…]
{refsection}
[references_supp.bib]
Supplementary Note 1 UNOSAT Dataset
Our reference dataset has been created by combining all existing UNOSAT datasets for Ukraine [unosat]. We combined them into 18 distinct areas of interest (AOIs). For our purposes, we only use the two strongest labels, destroyed and severely damaged. For the sake of completeness, we also report the third one here, moderately damaged. For Mariupol and Kharkiv, for which multiple analyses were conducted by UNOSAT, we keep only the latest label assigned to a point, assuming that in the case of repeated assessments, the annotations were progressively refined. Table S4 summarizes our AOIs and the number of labeled locations in each of them.
@@ Acknowledgments
Table S4: Overview of ground truth Summary of the 18 distinct areas of interest (AOI) in Ukraine, merging data from 29 UNOSAT products. The table lists counts of destroyed, severely damaged, and moderately damaged labels per area, totaling 16,548 labeled damage instances, or 10,934 after excluding the moderately damaged ones.
Supplementary Note 2 Detailed Comparison with PWTT
The Pixel-Wise T-Test (PWTT), proposed in [Ballinger2024], is a statistical approach designed to identify changes by comparing the mean values of two Sentinel-1 amplitude time series, adjusted by their standard deviation. We use it as a performance baseline since it utilizes the same data and has similar specifications, in particular, the scalability to large regions and the ease of deployment on GEE.
@@
For a meaningful comparison, we keep the recommended, optimal cutoff value but rerun PWTT using our time periods and our own, stricter definition of negative labels. We report the corresponding results in Table S5. For readability, we repeat our results in Table S6. It should be noted that Chernihiv and Kharkiv, which represent 61% of our test set, were used by the original authors during the calibration of the PWTT. Nevertheless, our method consistently outperforms the baseline across all metrics.
@@
Table S5: Evaluation metrics for PWTT. Quantitative performance of the PWTT approach on our test set, with the cutoff value of 1.63 as recommended for Ukraine. All metrics are given as percentages.
@@
Table S6: Evaluation metrics for our approach. Quantitative performance of our approach on our test set, with confidence threshold 0.5 (reprinted from Table 1). All metrics are given as percentages.
Supplementary Note 3 Choice of threshold
@@
There is an inherent trade-off between precision and recall, and the optimal threshold depends on the specific application. To keep false alarms at a low enough rate, we set a target precision of 90%, for which we find an optimal confidence threshold of 0.655. That value has been used to generate all metrics in the present paper unless stated otherwise. With this threshold, the user can expect a recall of 46.5%. In practice, users of our tool can dynamically adjust the confidence threshold to suit their specific needs. For instance, one could start with a high threshold to discover hotpots with a very high likelihood of destruction, then gradually decrease the threshold to increase the recall and better assess the extent of destruction. Fig. S6 illustrates how our metrics vary with different thresholds.
@@
Figure S6: Performance of the model for different thresholds. Pixel-level performance of our method on our test set, for different confidence thresholds. We set a target desired precision of 90% and found a confidence threshold of 0.655. The two vertical lines indicate the thresholds 0.5 and 0.655.
Supplementary Note 4 Temporal Evolution of Damage
@@
The series of maps in Figure Fig. S7 illustrates the progression of building damage over the first two years of the conflict. The maps clearly delineate the primary locations of battles and the frontline, highlighting the areas most affected as the conflict persists.
@@
Figure S7: Evolution of damage distribution in Ukraine over time. Percentage of buildings likely damaged for each period, cumulated since the beginning of the war and aggregated by hromadas. The predictions were thresholded at 0.655, and only buildings larger than 50 m2 were considered.
Supplementary Note 5 Limitations of Building Footprints
@@
The building footprints provided by Overture Maps [OvertureMaps], which for Ukraine were compiled from OpenStreetMap [OpenStreetMap] and Microsoft [MSFTfootprints], inevitably come with limitations. For instance, OpenStreetMap data may be obsolete due to the demolition of older structures (Fig. S8.A) or the construction of new ones (Fig. S8.B). While Microsoft footprints serve as a valuable complement to OpenStreetMap data, they were automatically retrieved with a deep neural network from Bing Maps imagery, which may not always offer recent images, and suffers from gaps due to cloudy or outdated acquisitions, for example in Kherson (Fig. S8.C). Moreover, it is not uncommon for large non-building structures, e.g. containers (Fig. S8.D), boats (Fig. S8.E) or even aircraft (Fig. S8.F), to be wrongly identified as buildings.
@@
Figure S8: Quality issues of existing building footprints. All images cover 500$\times$500 m2 and are displayed alongside the corresponding building footprints sourced from Overture Maps. Only buildings larger than 50 m2 are depicted. (A) Outdated buildings in Sevastopol. (B) Absent shopping mall near Chernihiv. (C) Missing footprints in Kherson. (D) Containers in Odessa. (E) Boats in Zaporizhzhia. (F) Antonov International Airport, Hostomel. Satellite images were obtained from ESRI/Maxar Technologies (A and C), Bing/Maxar Technologies (B, E, and F), and Google/Maxar Technologies (D).
Supplementary Note 6 World Bank Estimates
@@
In this section, we compare our estimates, with a total of $\approx$400,000 buildings damaged between Feb 2022 to Feb 2024, to figures provided by the World Bank (WB) in their third Rapid Damage and Needs Assessment report [WorldBank2023, p. 74-80].
@@
Fig. S9 shows the differences between our estimates and those of the WB. Excluding the other oblasts outside the direct conflict zone, both distributions align quite well, with indeed the majority of damage concentrated in Donetska and Kharkivska. However, our model estimates considerably more damage than the WB for the other oblasts, i.e., and the ones further from the frontline. It appears that construction and other non-conflict-related changes dominate in those regions, which our model is unable to distinguish from building damages. This highlights that there is a trade-off between, on the one hand, comprehensive screening to detect also unexpected conflict-induced destruction far from the known frontline; and, on the other hand, accurate assessment by ruling out areas that are a-priori unlikely to be affected.
@@
Figure S9: World Bank estimates vs Ours. Comparison of the geographical distribution of the World Bank estimates against ours. The World Bank provides spatial distribution for their damage cost estimates, while our spatial distribution shows buildings likely damaged.
Supplementary Note 7 Changes in Backscatter due to Destruction
@@
The backscatter signal from a building is influenced by numerous variables, including building characteristics (e.g., geometry, orientation, surface texture, material composition), satellite parameters (e.g., frequency, polarization, incidence angle, orbit direction), and environmental factors (e.g., vegetation, soil moisture, snow, or heavy rainfall). Intuitively, one might expect that intact structures yield stronger specular returns, with flat surfaces and right-angle corners acting as corner reflectors, whereas structural damage would disrupt these patterns, producing a more diffuse signal. However, several studies indicate that damage signatures in SAR data are highly variable and depend on the type and extent of destruction [Plank2014, Ge2020]. Damaged buildings can even exhibit backscatter signals as high as undamaged structures, or even higher [Brunner2010].
@@
Our experiments corroborate these findings. Fig. S10 presents pixel-wise backscatter time series that illustrate how conflict-related destruction appears in the data. High-resolution pre- and post-destruction satellite imagery from Google Earth is provided for reference. We observe that the backscatter amplitude following destruction can vary widely – increasing, decreasing, or remaining largely stable – depending on the location, the satellite orbit, and the polarization. These examples emphasize the need to utilize all available orbits in order to maximize reliability.
@@
Figure S10: Example of Sentinel-1 time series Sentinel-1 backscatter time series at selected UNOSAT locations for VV and VH polarizations, with VHR satellite images from Google Earth showing pre- and post-destruction conditions. Vertical dashed lines indicate the date of Russia’s full-scale invasion (black) and the dates of the displayed pre-destruction (green) and post-destruction (red) images. For readability, only two orbits per location are shown, even if in some locations up to four orbits are available. Amplitude can decrease (B, C), increase (A, C) or remain relatively stable (B, D) after destruction. The same event may manifest differently depending on the flight direction (B), or even on the incidence angle within the same flight direction (C). Basemaps from Google Earth/Maxar Technologies.
Supplementary Note 8 Ablation studies
@@
The $F_{1}$-score of our optimal configuration, which utilizes both VV and VH Sentinel-1 bands, all seven features, 50 decision trees and extracts the time series pixel-wise, is 0.72. Fig. S11 summarized the outcome for different settings. We observe the following characteristics:
@@
Input data: We tried replacing Sentinel-1 SAR data with Sentinel-2 optical images [sentinel_2]. Pixel-wise Sentinel-2 time series were extracted from the Level-2A collection available on GEE [s2Harmonized], from which we filter out cloudy pixels using GEE’s CloudScore+ mask [Pasquarella2023]. Various band combinations were tried, but incorporating Sentinel-2 data never enhanced model performance. This was expected, as our pipeline was design for Sentinel-1 properties, such as stable pre-event signals and a focus on temporal context over spatial context. Intuitively, it is harder to distinguish intact buildings from damaged ones by looking at the brightness change at an individual pixel location. Also combining Sentinel-1 and Sentinel-2 led to poorer results than using Sentinel-1 alone, likely due to the confounding effect of a six-fold increase in features, most of which contain little information for the task.
Model: We evaluated alternative classifiers available in GEE, i.e., Boosted Trees [Li2014Smile] and SVM [libvsm]. Boosted trees showed weaker performance, while taking approximately 3$\times$ longer to run. We were unable to obtain results for the SVM: when applied to our data it quickly exceeded the computational budget allowed by GEE.